# SMReferences for Appendix

### v3

@@ 5 Concluding remarks
In this paper, we present a novel methodology called Double-Calibration, which produces $\sqrt{n}$-consistent and asymptotic normal estimators for ATE and regression coefficient in high-dimensional semiparametric partially linear models under the (minimal sparsity) condition. There are several problems that are worth considering in future works. First, as mentioned in the end of the Introduction, we only construct $\sqrt{n}$-consistent and asymptotic normal estimators for ATE. Semiparametric efficient estimator under the (minimal sparsity) condition remains an open question. Second, here we only considered sparse GLMs. It is an open question to bridge the following four regimes in high-dimensional statistics: the sparse regime, the dense but $1\ll p\ll n$ regime (Liu et al., 2017; Liu and Li, 2023; Su et al., 2023), the dense but proportional asymptotic regime ($p/n\rightarrow c\in(0,\infty)$) (Yadlowsky, 2022; Jiang et al., 2022), and beyond ($p\gg n$) (Robins et al., 2008; Liu et al., 2021). Verzelen and Gassiat (2018) have made some attempt at this problem for quadratic functionals. Finally, we point out another interesting observation that warrants further study. When the nuisance parameters are modeled as sparse high-dimensional GLMs, at least for ATE, we have seen that $\sqrt{n}$-consistent estimation can be achieved by the proposed Double-Calibration strategy. However, in the context of classical nonparametric-type models (e.g. Hölder class) for the nuisance parameters, the only known $\sqrt{n}$-consistent estimators for ATE under minimal condition is based on higher-order influence functions (Robins et al., 2008; Liu et al., 2017; Liu and Li, 2023). It remains to be seen if similar idea to that developed in this paper can also be applied to the nonparametric setting.
Appendix A Proof of the main results
@@ A.1 Proof of Lemma 1
Let $W_{i}\coloneqq T_{i}/\pi_{i}^{*}-1$. Then by conditioning on $\{\mathbf{X},\mathcal{D}_{\mathrm{aux}},\mathcal{D}_{\mathrm{tr}}\}$, the $W_{i}$’s are independent sub-Gaussian random variables with variance proxy bounded below by a constant $(1-c_{\pi})/c_{\pi}$. From this, as a direct consequence of the Chernoff bound, we have by conditioning on $\{\mathbf{X},\mathcal{D}_{\mathrm{aux}},\mathcal{D}_{\mathrm{tr}}\}$, with probability converging to $1$, the constraints in (8) are satisfied with $\pi_{i}$ replaced by $\pi_{i}^{*}$ and by choosing $\eta_{\pi}\asymp\sqrt{\log p/n}$ sufficiently large.
From Condition 2, we have that all the $\pi_{i}^{*}$’s are in the domain space of the link function $\phi(\cdot)$, which means that the vector $\phi^{-1}({\bm{\pi}}^{*})$ is well defined. Moreover, from the construction of $\tilde{\mathbf{X}}$ and Condition 5, it is immediate that the matrix $\tilde{\mathbf{X}}$ is almost surely full row rank, so that there is at least one $\gamma$ such that almost surely the following equality holds,
@@
This proves the desired result. ∎
@@ A.2 Proof of Lemma 2
$\hat{\mu}_{\texttt{ora},i}\coloneqq\frac{\pi_{i}^{*}(r_{i}^{*}-\hat{r}_{i})}{\hat{\pi}_{i}}.$
@@
Under the same setup as Lemma 2, we have that with probability converging to $1$, $\hat{\bm{\mu}}_{\texttt{ora}}$ is a feasible solution to (5).
@@
This follows directly from the proof of Wang and Shah [2020, Theorem 19]. ∎
@@
Suppose that we are under the event where Lemma 8 holds true, then it is easy to see that with probability converging to $1$,
$\|\hat{\bm{\mu}}\|_{2}\leq\|\hat{\bm{\mu}}_{\texttt{ora}}\|_{2}\lesssim\sqrt{\sum_{i=1}^{n}(X_{i}^{\top}(\hat{\beta}-\beta^{*}))^{2}},$
where the last inequality follows from (i) the definition of $\hat{\pi}$ and (ii) that $\psi^{\prime}(\cdot)$ is uniformly bounded since we are under Condition 7. Now applying a Bernstein’s inequality [Vershynin, 2018, Lemma 2.7.7], with probability converging to $1$, we have
$\|\hat{\bm{\mu}}\|_{2}\lesssim\sqrt{n}\|\hat{\beta}-\beta^{*}\|_{2}.$
The desired result is then a direct consequence of Condition 7 and the sparsity constraint $s_{r}\ll\sqrt{n}/\log p$. ∎
@@ A.3 Proof of Theorem 3
First, we have that for some $\eta_{i}\in[0,1]$ for $i=1,\cdots,n$, the following decomposition holds
$\hat{\tau}_{\textrm{DCal}}$ $-\bar{\tau}^{\ast}$
$=$ $\ \underset{\mathrm{I}}{\underbrace{\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}}{\tilde{\pi}_{i}}-1\right)\psi^{\prime}(X_{i}^{\top}\hat{\beta})X_{i}^{\top}(\hat{\beta}-\beta^{*})}}$
$+\underset{\mathrm{II}}{\underbrace{\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}}{\tilde{\pi}_{i}}-1\right)\psi^{\prime\prime}(\eta_{i}X_{i}^{\top}\hat{\beta}+(1-\eta_{i})X_{i}^{\top}\beta^{*})\left(X_{i}^{\top}(\hat{\beta}-\beta^{*})\right)^{2}}}$
$+\underset{\mathrm{III}}{\underbrace{\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}}{\tilde{\pi}_{i}}-1\right)\hat{\mu}_{i}}}+\frac{1}{n}\sum_{i=1}^{n}\frac{T_{i}\varepsilon_{i}(1)}{\tilde{\pi}_{i}}.$
As explained in Remark 2, the last term in the above decomposition has mean zero by design of the program (8). Invoking constraint (9) and Condition 7, we have that with probability converging to $1$,
@@
where for the last inequality we use the assumption that $\psi^{\prime}$ is uniformly bounded since we are under Condition 7. Under Condition 5, we further have that almost surely, $\max_{j}\|\mathbf{X}_{j}\|_{2}/\sqrt{n}\lesssim 1$. Putting these together and recalling that $s_{r}=o(\sqrt{n}/\log p)$, we can conclude that $|\mathrm{I}|=o_{\mathbb{P}}(1)$.
For $\mathrm{II}$, since Conditions 6 and 7 hold, together with constraint (12), we have that, with probability converging to $1$,
@@
In light of this upper bound and the Bernstein’s inequality [Vershynin, 2018, Lemma 2.7.7], with probability converging to $1$, we have that $\mathrm{II}\lesssim\|\hat{\beta}-\beta^{*}\|_{2}^{2}$. Then it immediately follows that $\mathrm{II}=o_{\mathbb{P}}(1)$ using again Condition 7 and that $s_{r}=o(\sqrt{n}/\log p)$.
For $\mathrm{III}$, under constraint (10) and the conclusion of Lemma 2, we have that with probability converging to $1$,
$|\mathrm{III}|\lesssim\sqrt{s_{r}}\frac{\log p}{n}\lesssim s_{r}\frac{\log p}{n}.$
Given the above control of $\mathrm{I}$, $\mathrm{II}$ and $\mathrm{III}$, the desired result follows. ∎
@@ A.4 Proof of Lemma 4
From Lemma 1, we have that with probability converging to $1$, $\gamma^{*}$ is a feasible solution to (8), so that
$\|\tilde{\gamma}-\hat{\gamma}\|_{1}\lesssim\|\gamma^{*}-\hat{\gamma}\|_{1}\lesssim s_{\pi}\sqrt{\frac{\log p}{n}},$
where for the last inequality we invoke Condition 8. Moreover,
$\|\tilde{\gamma}-\gamma^{*}\|_{1}\leq\|\tilde{\gamma}-\hat{\gamma}\|_{1}+\|\hat{\gamma}-\gamma^{*}\|_{1}\lesssim s_{\pi}\sqrt{\frac{\log p}{n}}.$
We now focus on proving the second result. Using that we are under constraint (11), we have that
$\left\|\frac{1}{n}\sum_{i=1}^{n}(T_{i}/\tilde{\pi}_{i}-1)\tilde{X}_{i}^{\top}\right\|_{\infty}\lesssim\sqrt{\frac{\log p\vee n}{n}}.$
Using the same analysis as in Lemma 1, we have that with probability converging to $1$, the above inequality holds, but with $\tilde{\pi}_{i}$ replaced by $\pi_{i}^{*}$. Putting these together yields
$\left\|\frac{1}{n}\sum_{i=1}^{n}(T_{i}/\tilde{\pi}_{i}-T_{i}/\pi_{i}^{*})\tilde{X}_{i}^{\top}\right\|_{\infty}\lesssim\sqrt{\frac{\log p\vee n}{n}}.$
From above, and by applying Hölder’s inequality, we have that
$\left|\frac{1}{n}\sum_{i=1}^{n}(T_{i}/\tilde{\pi}_{i}-T_{i}/\pi_{i}^{*})\tilde{X}_{i}^{\top}(\tilde{\gamma}-\gamma^{*})\right|\lesssim s_{\pi}\frac{\sqrt{\log p\log p\vee n}}{n}.$
@@
From Condition 5, we have that for all $i$, $|X_{i}^{\top}(\tilde{\gamma}-\gamma^{*})|\leq\|X_{i}\|_{\infty}\|\tilde{\gamma}-\gamma^{*}\|_{1}\lesssim 1$. Using Condition 2, we have that $|X_{i}^{\top}\gamma^{*}|\lesssim 1$ as well. From this, and the lower bound in Condition 9, we have that
@@
Now in light of the constraint (12), we further have with probability converging to $1$,
@@
Finally, applying Bernstein’s inequality and Condition 2, we have that with probability converging to $1$,
@@
$\lesssim s_{\pi}\frac{\sqrt{\log p\log p\vee n}}{n}+\left\|\frac{\tilde{\mathbf{X}}^{\top}\tilde{\mathbf{X}}}{n}-\frac{1}{n}\sum_{i=1}^{n}\frac{T_{i}}{\pi_{i}^{*}}\tilde{X}_{i}\tilde{X}_{i}^{\top}\right\|_{\infty}\|\tilde{\gamma}-\gamma^{*}\|_{1}^{2}.$
By Hoeffding’s inequality, with probability converging to $1$,
@@
From above, we conclude that

### v4

@@ 6 Concluding remarks
In this paper, we present a novel methodology called Double-Calibration, which produces $\sqrt{n}$-consistent and asymptotic normal estimators for ATE and regression coefficient in high-dimensional semiparametric partially linear models under the (minimal sparsity) condition. There are several problems that are worth considering in future works. First, as mentioned in the end of the Introduction, we only construct $\sqrt{n}$-consistent and asymptotic normal estimators for ATE. Semiparametric efficiency bound following Le Cam’s approach under the (minimal sparsity) condition remains an open question, but ideas in Mukherjee and Sen (2020) might be promising directions to explore. Second, here we only considered high-dimensional sparse GLMs. It is an open question to bridge the following five regimes in high-dimensional statistics: the sparse but $p\gg n$ regime (our case), the sparse but $p\leq n$ regime, the dense but $1\ll p\ll n$ regime (Liu et al., 2017; Liu and Li, 2023; Su et al., 2023), the dense but proportional asymptotic regime ($p/n\rightarrow c\in(0,\infty)$) (Yadlowsky, 2022; Jiang et al., 2022; Celentano and Wainwright, 2023; Chen et al., 2024), and beyond ($p\gg n$ but no sparsity) (Robins et al., 2008, 2023; Liu et al., 2021). Verzelen and Gassiat (2018) and Kong and Valiant (2018) have made some attempt at this problem for quadratic functionals. Third, our framework might also be extended to more complex causal inference settings, such as high-dimensional mediation analysis (Lin et al., 2023) or high-dimensional proximal causal learning (Deaner, 2021). Finally, we point out another interesting observation that warrants further study. When the nuisance parameters are modeled as sparse high-dimensional GLMs, at least for ATE, we have seen that $\sqrt{n}$-consistent estimation can be achieved by the proposed Double-Calibration strategy. However, in the context of classical nonparametric-type models (e.g. Hölder class) for the nuisance parameters, the only known $\sqrt{n}$-consistent estimators for ATE under minimal condition is based on higher-order influence functions (Robins et al., 2008; Liu et al., 2017; Liu and Li, 2023). It remains to be seen if similar idea to that developed in this paper can also be applied to the nonparametric setting or vice versa.
Acknowledgments
The authors are grateful for the extremely insightful discussions and comments by Rajarshi Mukherjee and Jamie Robins. The simulations in this paper were run on the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University. LL is affiliated with the Shanghai Artificial Intelligence Laboratory and The Smart Justice Lab at the Koguan Law School of Shanghai Jiao Tong University. LL gratefully acknowledges funding support by NSFC Grant No.12471274 and No.12090024, Shanghai Municipal Science and Technology Grant No.21ZR1431000, No.2021SHZDZX0102, and No.21JC1402900.
Appendix A Proofs of the results concerning Section 3
@@ A.1 Proof of Lemma 1
Let $W_{i}\coloneqq T_{i}/\pi_{i}^{*}-1$. Then by conditioning on $\{\tilde{\mathbf{X}},\hat{\bm{\mu}},\hat{\gamma},\hat{\beta}\}$, the $W_{i}$’s are independent sub-Gaussian random variables with variance proxy bounded below by a constant $(1-c_{\pi})/c_{\pi}$. From this, as a direct consequence of the Chernoff bound and the fact that almost surely, $T_{i}/\pi_{i}^{*}\leq c_{\pi}^{-1}$ since we are under Condition 2, there exist constants $c_{1},c_{2},c^{\prime}>0$ and $M_{\pi}\geq c_{\pi}^{-1}$ sufficiently large such that by choosing $\eta_{\pi_{1}}\leq c_{1}\sqrt{\frac{\log p}{n}},\eta_{\pi_{2}}\leq c_{2}\sqrt{\frac{\log p\vee n}{n}}$ for any realization of $\{\tilde{\mathbf{X}},\hat{\bm{\mu}},\hat{\gamma},\hat{\beta}\}$,
[…]
Second, from Condition 2, we have that all the $\pi_{i}^{*}$’s are in the domain space of the link function $\phi(\cdot)$, which means that the vector $\phi^{-1}({\bm{\pi}}^{*}):=(\phi^{-1}(\pi^{*}_{1}),\ldots,\phi^{-1}(\pi^{*}_{n}))^{\top}$ is well defined. Moreover, from the construction of $\tilde{\mathbf{X}}$ and Condition 4(ii), it is immediate that the matrix $\tilde{\mathbf{X}}$ is almost surely full row rank, so that almost surely there is at least one $\gamma$ satisfying
@@
Putting together, we prove the desired result. ∎
@@ A.2 Proof of Lemma 2
$\hat{\mu}_{\texttt{ora},i}\coloneqq\frac{\pi_{i}^{*}(r_{i}^{*}-\hat{r}_{i})}{\hat{\pi}_{i}},\hat{\bm{\mu}}_{\texttt{ora}}\coloneqq(\hat{\mu}_{\texttt{ora},1},\cdots,\hat{\mu}_{\texttt{ora},n})^{\top}.$
@@
Suppose Conditions 12 and 45 hold and let the constant $M_{\gamma}$ be given; then by choosing $\eta_{r}\asymp\sqrt{\frac{\log p}{n}},M_{r}\asymp 1$ large enough, with probability converging to $1$, $\hat{\bm{\mu}}_{\texttt{ora}}$ is a feasible solution to (6b)–(6c).
@@
Suppose $\mathcal{D}_{\mathrm{tr}}$ is under the event defined by Condition 5, we have that by conditioning on $\mathcal{D}_{\mathrm{tr}}$, the $j$-th index of the vector ${\mathbf{X}}_{\mathrm{aux}}^{\top}{\mathbf{\Pi}}_{\mathrm{aux}}\tilde{\mathbf{R}}_{\mathrm{aux}}$ is a sum of $n$ i.i.d. sub-exponential random variables, so that by Bernstein’s inequality and a union bound, we have that given any $m\in\mathbb{N}$, there exist positive constants $c_{1}$$c_{4}$ depending on $m$ such that when $\mathcal{D}_{\mathrm{tr}}$ is under the event defined by Condition 5,
$\mathbb{P}\left(\left\|\frac{1}{n_{\mathrm{aux}}}{\mathbf{X}}_{\mathrm{aux}}^{\top}{\bm{\Pi}}_{\mathrm{aux}}\tilde{\mathbf{R}}_{\mathrm{aux}}-\mathbb{E}\left[\frac{1}{n_{\mathrm{aux}}}{\mathbf{X}}_{\mathrm{aux}}^{\top}{\bm{\Pi}}_{\mathrm{aux}}\tilde{\mathbf{R}}_{\mathrm{aux}}\mid\mathcal{D}_{\mathrm{tr}}\right]\right\|_{\infty}\geq c_{1}\sqrt{\frac{\log p}{n_{\mathrm{aux}}}}\mid\mathcal{D}_{\mathrm{tr}}\right)\leq c_{2}p^{-m},$
and
$\mathbb{P}\left(\left\|\frac{1}{n}{\mathbf{X}}^{\top}{\bm{\Pi}}\hat{\bm{\mu}}_{\texttt{ora}}-\mathbb{E}\left[\frac{1}{n}{\mathbf{X}}^{\top}{\bm{\Pi}}\hat{\bm{\mu}}_{\texttt{ora}}\mid\mathcal{D}_{\mathrm{tr}}\right]\right\|_{\infty}\geq c_{3}\sqrt{\frac{\log p}{n}}\mid\mathcal{D}_{\mathrm{tr}}\right)\leq c_{4}p^{-m}.$
Since $\mathcal{D}$ and $\mathcal{D}_{\mathrm{aux}}$ are i.i.d. sampled from the same distribution, we can obtain that almost surely,
$\mathbb{E}\left[\frac{1}{n}{\mathbf{X}}^{\top}{\bm{\Pi}}\hat{\bm{\mu}}_{\texttt{ora}}\mid\mathcal{D}_{\mathrm{tr}}\right]=\mathbb{E}\left[\frac{1}{n_{\mathrm{aux}}}{\mathbf{X}}_{\mathrm{aux}}^{\top}{\bm{\Pi}}_{\mathrm{aux}}\tilde{\mathbf{R}}_{\mathrm{aux}}\mid\mathcal{D}_{\mathrm{tr}}\right].$
In light of the above three inequalities and a union bound, we prove that (6b) can be satisfied with probability converging to $1$ by choosing $\eta_{r}\asymp\sqrt{\frac{\log p}{n}}$ sufficiently large.
The control of (6c) is straightforward by simply setting $M_{r}:=\frac{m_{r}+m_{Y}}{\phi(-M_{\gamma})}$. ∎
@@
Suppose that we are under the event where Lemma 8 holds true, then it is easy to see that almost surely,
$\|\hat{\bm{\mu}}\|_{2}\leq\|\hat{\bm{\mu}}_{\texttt{ora}}\|_{2}\lesssim\sqrt{\sum_{i=1}^{n}(X_{i}^{\top}(\hat{\beta}-\beta^{*}))^{2}},$ (22)
where the last inequality follows from (i) the definition of $\hat{\pi}$ and (ii) that $\psi^{\prime}(\cdot)$ is uniformly bounded since we are under Condition 6.
Now applying a Bernstein’s inequality [Vershynin, 2018, Lemma 2.7.7], we have that there exists constants $c_{1},c_{2}>0$ such that almost surely,
$\mathbb{P}\left(\|\mathbf{X}(\hat{\beta}-\beta^{*})\|_{2}/\sqrt{n}\geq c_{1}\|\hat{\beta}-\beta^{*}\|_{2}\mid\hat{\beta}\right)\leq e^{-c_{2}n}.$ (23)
In light of both inequalities, the desired result follows from $s_{r}=o(\sqrt{n}/\log p)$. ∎
@@ A.3 Proof of Theorem 3
Using Taylor expansion, we have that for some $\eta_{i}\in[0,1]$ for $i=1,\cdots,n$, the following decomposition holds
$\begin{split}\hat{\tau}_{\mathrm{DCal}}&-\bar{\tau}^{\ast}\\ =&\ \underset{\mathrm{I}}{\underbrace{\frac{1}{n}\sum_{i=1}^{n}\left(1-\frac{T_{i}}{\tilde{\pi}_{i}}\right)\psi^{\prime}(X_{i}^{\top}\hat{\beta})X_{i}^{\top}(\hat{\beta}-\beta^{*})}}\\ &+\underset{\mathrm{II}}{\underbrace{\frac{1}{n}\sum_{i=1}^{n}\left(1-\frac{T_{i}}{\tilde{\pi}_{i}}\right)\psi^{\prime\prime}(\eta_{i}X_{i}^{\top}\hat{\beta}+(1-\eta_{i})X_{i}^{\top}\beta^{*})\left(X_{i}^{\top}(\hat{\beta}-\beta^{*})\right)^{2}}}\\ &+\underset{\mathrm{III}}{\underbrace{\frac{1}{n}\sum_{i=1}^{n}\left(1-\frac{T_{i}}{\tilde{\pi}_{i}}\right)\hat{\mu}_{i}}}+\frac{1}{n}\sum_{i=1}^{n}\frac{T_{i}\varepsilon_{i}(1)}{\tilde{\pi}_{i}}.\end{split}$ (24)
We first consider $\mathrm{I}$. Invoking Lemma 1, we have that the constraint (9b) holds with probability converging to $1$. In light of this and Condition 6, we have that with probability converging to $1$,
@@
where for the last inequality we use the assumption that $\psi^{\prime}$ is uniformly bounded since we are under Condition 6. Under Condition 4(ii), we further have that almost surely, $\max_{j}\|\mathbf{X}_{j}\|_{2}/\sqrt{n}\lesssim 1$. Putting these together and recalling that $s_{r}=o(\sqrt{n}/\log p)$, we can conclude that $|\mathrm{I}|=o_{\mathbb{P}}(n^{-1/2})$.
For $\mathrm{II}$, since Conditions 5 and 6 hold, together with constraint (9e), we have that, with probability converging to $1$,
@@
In light of this upper bound and recall (23), we have that with probability converging to $1$, we have that $\mathrm{II}\lesssim\|\hat{\beta}-\beta^{*}\|_{2}^{2}$. Then it immediately follows that $\mathrm{II}=o_{\mathbb{P}}(n^{-1/2})$ using again Condition 6 and that $s_{r}=o(\sqrt{n}/\log p)$.
For $\mathrm{III}$, as a direct consequence of (9c) and the conclusion of Lemma 2, we have that
$|\mathrm{III}|=o_{\mathbb{P}}\left(\frac{1}{\sqrt{n}}\cdot\sqrt{\frac{\log p}{\sqrt{n}}}\right)=o_{\mathbb{P}}(n^{-1/2}).$
Given the above controls of $\mathrm{I}$, $\mathrm{II}$ and $\mathrm{III}$, the desired result follows. ∎
@@ A.4 Proof of Lemma 4
Without loss of generality, throughout this section we assume that $p\geq n$, so that $\tilde{X}_{i}$ and $X_{i}$ are equivalent. The proof of the case $p<n$ can be obtained via applying an analogous argument. From Lemma 1, we have that with probability converging to $1$, $\gamma^{*}$ is a feasible solution to (9a) when the tuning parameters are chosen according to Lemma 1, so that with probability converging to $1$,
$\|\tilde{\gamma}-\hat{\gamma}\|_{1}\leq\|\gamma^{*}-\hat{\gamma}\|_{1}\lesssim s_{\pi}\sqrt{\frac{\log p}{n}},$
where for the last inequality we invoke Condition 7. Moreover,
$\|\tilde{\gamma}-\gamma^{*}\|_{1}\leq\|\tilde{\gamma}-\hat{\gamma}\|_{1}+\|\hat{\gamma}-\gamma^{*}\|_{1}\lesssim s_{\pi}\sqrt{\frac{\log p}{n}}.$ (25)
We now focus on proving the second result. Working under the event given by Lemma 1, using that we are under constraint (9d), we have that
$\left\|\frac{1}{n}\sum_{i=1}^{n}(T_{i}/\tilde{\pi}_{i}-1)\tilde{X}_{i}^{\top}\right\|_{\infty}\lesssim\sqrt{\frac{\log p\vee n}{n}},$
and that the above inequality still holds with $\tilde{\pi}_{i}$ replaced by $\pi_{i}^{*}$. Putting these together yields that under the event given by Lemma 1,
$\left\|\frac{1}{n}\sum_{i=1}^{n}(T_{i}/\tilde{\pi}_{i}-T_{i}/\pi_{i}^{*})\tilde{X}_{i}^{\top}\right\|_{\infty}\lesssim\sqrt{\frac{\log p\vee n}{n}}.$ (26)
In light of (25) and (26), and by applying Hölder’s inequality, we have that with probability converging to $1$,
$\left|\frac{1}{n}\sum_{i=1}^{n}(T_{i}/\tilde{\pi}_{i}-T_{i}/\pi_{i}^{*})\tilde{X}_{i}^{\top}(\tilde{\gamma}-\gamma^{*})\right|\leq\left\|\frac{1}{n}\sum_{i=1}^{n}(T_{i}/\tilde{\pi}_{i}-T_{i}/\pi_{i}^{*})\tilde{X}_{i}^{\top}\right\|_{\infty}\|\tilde{\gamma}-\gamma^{*}\|_{1}\lesssim s_{\pi}\frac{\sqrt{\log p\log p\vee n}}{n}.$
@@
From Conditions 4(ii) and (25), we have that with probability converging to $1$, for all $i$, $|X_{i}^{\top}(\tilde{\gamma}-\gamma^{*})|\leq\|X_{i}\|_{\infty}\|\tilde{\gamma}-\gamma^{*}\|_{1}\lesssim 1$. Using Condition 2 and the constraint on the link function provided by Condition 3(ii), we have that almost surely $|X_{i}^{\top}\gamma^{*}|\lesssim 1$ as well. From this, and the lower bound in Condition 8, we have that there exists a constant $c>0$ such that with probability converging to $1$, for all $i$,
$\phi^{\prime}(\iota_{i}X_{i}^{\top}\gamma^{*}+(1-\iota)X_{i}^{\top}\tilde{\gamma})\geq c,$
from which we have
@@
This, in combination with  (9e), yields
@@
Finally, recall that we have (23) and Condition 2, we have that with probability converging to $1$,
@@
$\lesssim s_{\pi}\frac{\sqrt{\log p\log p\vee n}}{n}+\left\|\frac{\tilde{\mathbf{X}}^{\top}\tilde{\mathbf{X}}}{n}-\frac{1}{n}\sum_{i=1}^{n}\frac{T_{i}}{\pi_{i}^{*}}\tilde{X}_{i}\tilde{X}_{i}^{\top}\right\|_{\infty}\|\tilde{\gamma}-\gamma^{*}\|_{1}^{2},$
where the for the last inequality we apply Hölder’s inequality. By Hoeffding’s inequality and a union bound, with probability converging to $1$,
@@
From above, we conclude that with probability converging to $1$,